### v1

# Probing LLMs for Multilingual Discourse Generalization Through a Unified Label Set

Abstract
Discourse understanding is essential for many NLP tasks, yet most existing work remains constrained by framework-dependent discourse representations. This work investigates whether large language models (LLMs) capture discourse knowledge that generalizes across languages and frameworks. We address this question along two dimensions: (1) developing a unified discourse relation label set to facilitate cross-lingual and cross-framework discourse analysis, and (2) probing LLMs to assess whether they encode generalizable discourse abstractions. Using multilingual discourse relation classification as a testbed, we examine a comprehensive set of $23$ LLMs of varying sizes and multilingual capabilities. Our results show that LLMs, especially those with multilingual training corpora, can generalize discourse information across languages and frameworks. Further layer-wise analyses reveal that language generalization at the discourse level is most salient in the intermediate layers. Lastly, our error analysis provides an account of challenging relation classes.
Probing LLMs for Multilingual Discourse Generalization
Through a Unified Label Set
Florian Eichinthanks: Equal contribution., Yang Janet Liu11footnotemark: 1, Barbara Plank, and Michael A. Hedderich MaiNLP, Center for Information and Language Processing, LMU Munich, Germany Munich Center for Machine Learning (MCML) {feichin,yliu,bplank,hedderich}@cis.lmu.de
1 Introduction
Many approaches to NLP primarily focus on sentence-level analyses (e.g. Heinzerling and Strube 2019; Pimentel et al. 2021; Mrini et al. 2020). However, there are many research questions which cannot be answered without considering sentences in a larger discourse: new meanings emerge from the relationships between sentences, and since more than one interpretation can be created, how do we determine the intended, most reasonable or justifiable meaning Schiffrin et al. (2015)?
Figure 1: Examples of the core discourse relation condition Bunt and Prasad (2016) annotated in different frameworks and languages using different labels.
Despite significant progress in discourse processing Webber et al. (2024); Zeldes et al. (2024); Stede (2011), much of the research and resources remain constrained by theory-/framework-specific assumptions, limiting the generalizability of findings across languages, domains, and communicative intents Liu and Zeldes (2023). This leads to datasets that are tightly coupled to their respective frameworks and limits the development of generalizable discourse models. While there has been work on framework-dependent parsing that leverages resources from other frameworks Braud et al. (2016) or languages Braud et al. (2017); Liu et al. (2021), the reliance on framework-specific corpora, which are typically scarce and skewed towards high-resource languages, further exacerbates the challenge of multilingual discourse processing. Thus, we need a unified view and approach to investigating discourse generalization.
From the theory and data perspectives, as argued in Bunt and Prasad (2016) and exemplified in Figure 1, despite differences between frameworks, there exists a set of ‘core’ discourse relations which are commonly found in existing approaches to discourse relations and their annotation. From the model perspective, there is growing evidence demonstrating that large language models (LLMs) learn and share generalizable abstraction across typologically diverse languages (e.g. Brinkmann et al. 2025; Peng and Søgaard 2024), but such capabilities remain underexplored in discourse.
In this work, we address discourse generalization across two dimensions using discourse relation classification as a testbed: we first develop a unified discourse relation label set to enable cross-lingual and cross-framework discourse analysis on the the multilingual DISRPT benchmark Braud et al. (2024). Then, we use probing (Alain and Bengio, 2018) to understand the internal discourse representations of $23$ LLMs with varying sizes and multilingual capabilities. We investigate whether their representations capture generalizable discourse abstractions across typologically diverse languages or whether they are limited by dataset-specific biases. We hypothesize that multilingual models might encode universal representations of certain relations while adjusting to language-specific features.
We find that overall LLMs are able to generalize at the discourse level across languages and frameworks, and that multilingual training and larger model sizes both increase probe performance. Our layer-wise analyses show that language generalization at the discourse level is most salient in the intermediate layers, which are most predictive of multilingual discourse information. To our best knowledge, this is the first work to apply a unified label set to multilingual discourse relation classification at scale. Lastly, we discuss challenges and biases in LLM discourse representations, providing insights into the limitations and potential avenues for improving discourse modeling in multilingual and generalization settings.
2 Related Work
Unifying Discourse Relations.
Discourse relations are fundamental to structuring coherent text and conveying meaning beyond the sentence level. Being able to identify and interpret these relations is crucial for many downstream NLP tasks, including machine comprehension Narasimhan and Barzilay (2015); Li et al. (2020), sentiment analysis Huber and Carenini (2019), question answering Chai and Jin (2004), and summarization Durrett et al. (2016); Cohan et al. (2018); Xu et al. (2020); Adams et al. (2023). However, due to different approaches to discourse relations, such as the Rhetorical Structure Theory (RST, Mann and Thompson 1988), the Penn Discourse Treenbank (PDTB, Webber et al. 2019), the Segmented Discourse Representation Theory (SDRT, Asher and Lascarides 2003), Discourse Dependency Structure (DEP, Li et al. 2014; Morey et al. 2018), and the Cognitive Approach to Coherence Relations (CCR, Sanders et al. 1992), researchers have not reached a consensus on a unified set of discourse relations. There have been a few mapping proposals and examinations on existing annotations Chiarcos (2014); Benamara and Taboada (2015); Bunt and Prasad (2016); Rehbein et al. (2016); Sanders et al. (2021); Demberg et al. (2019), but they are either merely focused on two frameworks at a time (e.g. RST and SDRT in Benamara and Taboada 2015), or on high-resource languages and news-centric data such as mapping RST-DT and PDTB v2 by Demberg et al. (2019) and PDTB v3 by Costa et al. (2023).
A similar effort to unify discourse relations has been made by the 2021 DISRPT Shared Task Zeldes et al. (2021) which introduced the first iteration of the discourse relation classification task in a unified format. It leveraged shared foundational assumptions across frameworks. However, no unified discourse relation labels were proposed, meaning that each dataset has its own label set, even for the ones that come from the same framework. To address this, we first conduct an extensive review of previous work on discourse relation mapping proposals and build upon existing work to present a unified label set, which we describe in Section 3.
Probing for Linguistic Representation and Generalization.
A growing body of research has explored the extent to which pretrained language models (PLMs) and LLMs encode linguistic representations and exhibit generalizable abstraction Hupkes et al. (2023). They primarily focus on probing morphology Brinkmann et al. (2025), syntax Conneau et al. (2018); Hale and Stanojević (2024), semantic knowledge Jumelet et al. (2021), and syntax-semantics comprehension through cognitive linguistics paradigms such as construction grammar Weissweiler et al. (2022). While some studies demonstrate that LLMs share latent grammatical representations across diverse languages Brinkmann et al. (2025), others also highlight key limitations in the semantic capabilities of LLMs Scivetti et al. (2025).
Previous work has examined the ability of PLMs/LLMs to understand discourse such as coreference resolution Gan et al. (2024); Saputa et al. (2024), but their investigations are either limited to framework-dependent representations, monolingual datasets, or focus on single domains. Specifically, Koto et al. (2021) examined a variety of PLMs through a set of framework-dependent probing tasks for discourse coherence by looking at the residual stream (i.e. token representations), while we approach discourse relation classification with a unified format and label set across various frameworks and $13$ languages (§3) using attentions (§4), offering opportunities for investigating discourse generalization across languages and frameworks. Kurfalı and Östling (2021) extended discourse probing to multilingual PLMs such as the multilingual BERT and XLM-RoBERTa (Devlin et al., 2019; Conneau et al., 2020) to examine how well they transfer discourse-level knowledge across languages, but their evaluation of discourse coherence was also framework-dependent and was only performed on two English datasets in the news domain. Lastly, Kim and Schuster (2023) studied discourse understanding in LMs by probing their ability to track discourse entities, but their investigation is also limited to English.
3 A Unified Label Set
Building up on previous work and effort on mapping discourse relations across corpora and frameworks Benamara and Taboada (2015); Bunt and Prasad (2016); Liu and Zeldes (2023), we present a unified set of $17$ discourse relation classes that are identified as core discourse relations by Bunt and Prasad (2016) and are attested in various domains and text types. This unified label set is motivated by both theoretical groundings and empirical studies, and takes annotation guidelines into considerations. Specifically, the proposed unified label set adapts the top-level classes from the mapping proposal described in Benamara and Taboada (2015) and extends it to phenomena frequent in dialogues such as acknowledgment, interruption, and correction Asher et al. (2016). We describe the top-level classes below and include definitions and examples for all $17$ relation labels in Appendix A.
temporal is mapped to framework-specific labels that establish a chronological sequence between events or states. Temporal relations indicate when one event occurs in relation to another such as before, after, or simultaneously. These relations help organize discourse by providing a timeline of events. RST’s sequence, PDTB’s Temporal.Asynchronous/Synchronous, and SDRT’s temploc and flashback (following Muller et al. (2012)) all fall under this class.
structuring corresponds to fine-grained discourse relations that organize the structure of a text or conversation without necessarily conveying content-based meaning, connect discourse units of distinct context and equal prominence, and help guide the reader or listener through the discourse. RST-style relations such as list and textual-organization, PDTB’s expansion.disjunction, and parallel and alternation in SDRT are mapped to this class.
thematic is a broad class which includes relations among the content of the propositions, according to Benamara and Taboada (2015). We adapt this top-level class to contain six subclasses: framing, attribution, mode, reformulation, comparison, and elaboration. In particular, we introduce reformulation, which corresponds to relations by which one discourse unit re-expresses the meaning of another in a different form and/or from a different perspective to help reinforce understanding. RST’s summary and restatement and PDTB’s expansion.equivalence are mapped to reformulation.
causal-argumentative contains subclasses that indicate a causal relation or involve rhetorical reasoning that shapes the coherence and persuasiveness of an argument: causal, adversative, explanation, evaluation, contingency, and enablement. adversative is defined as connecting discourse units for which some incompatibility is being highlighted and covers commonly used discourse relations in virtually all frameworks such as concession and contrast.
topic-management contains three subclasses that cover discourse phenomena that involve interaction, topic shifts, and question-answer pairs: topic-adjustment, topic-change, and topic-comment. topic-adjustment is primarily used for cases where a discourse segment modifies, redirects, or adjusts the ongoing topic of discussion such as interruption, which signal deviations from the expected discourse progression, often reflecting interactive or dynamic aspects of communication.
4 Probing LLMs for Discourse Relations
Discourse relation classification is the task of identifying the coherence relations that hold between different parts of a text, such as recognizing that one sentence specifies the cause of events in another, or that a subordinate clause indicates the condition of a main clause Jurafsky and Martin (2025), as exemplified in Figure 1. Successfully solving this task requires combining the information contained in different parts of these sequences. We aim to investigate whether current LLMs have access to and process this information. Since almost all current state-of-the-art open-source models are decoder-only Transformer models (Vaswani et al., 2017), we focus on this particular architecture.
In the computational graph of the Transformer, attention layers provide the only connections between the next token prediction and previous token positions of the input sequence. We therefore hypothesize that generating predictions relying on discourse-level information necessarily has to involve the connections provided by the self-attention layers. To uncover the extent to which discourse information is represented, we thus propose to probe the attention scores between the tokens contained in the two input sequences for discourse relation classification.
Probing requires a fixed length representation to enable training a classifier. Inspired by Alain and Bengio (2018) and Koto et al. (2021), we use maximum pooling to turn the attention scores of variable token-length sentences into a fixed-length relational representation. To be more precise, for a document $d$ of length $N$ tokens, we propose to compute the full attention matrix $\mathbf{X}\in\mathbb{R}^{L\times H\times N\times N}$ where $L$ is the number of layers and $H$ is the number of attention heads in each layer. Let $I_{1}=(i_{1},i_{2},...,i_{N_{1}})$, $I_{2}=(j_{1},j_{2},...,j_{N_{2}})$ be the token indices of two sequences $s_{1},s_{2}\subset d$ where $I_{1}<I_{2}$. The pooling step is carried out for each attention head in each layer. We thus have:
$\begin{split}\mathbf{C}&=\left(\max(\{\mathbf{X}_{i,j,k,l}|k\in I_{1},l\in I_{2}\})\right)_{i,j}\\ \mathbf{D}_{m}&=\left(\max(\{\mathbf{X}_{i,j,k,l}|k\in I_{m},l\in I_{m}\})\right)_{i,j}\\ \end{split}$ (1)
In other words, $\mathbf{C},\mathbf{D}_{1},\mathbf{D}_{2}$ are the matrices of pooled attention scores between and within the two spans respectively (we only consider the lower half of the attention matrix, as for decoder models the upper half is masked). We probe on the flattened concatenation $\mathbf{a}\in\mathbb{R}^{3LH}$ of these representations. Here, we follow Tenney et al. (2019) and use a two-layer MLP with $\mathrm{tanh}$ and Sigmoid activations $\sigma$. Our probe is thus a classifier of the form
$\mathbf{y}=\sigma(\mathbf{W}_{2}\mathrm{tanh}(\mathbf{W}_{1}\mathbf{a}+\mathbf{b}_{1})+\mathbf{b}_{2})$ (2)
where $\mathbf{W}_{1}\in\mathbb{R}^{D\times 3LH},\mathbf{W}_{2}\in\mathbb{R}^{C\times D},\mathbf{b}_{1}\in\mathbb{R}^{D},\mathbf{b}_{2}\in\mathbb{R}^{C}$ are parameters and $\mathbf{y}$ is the one-hot encoded target vector. The hidden size $D$ is a hyperparameter (more training details in Appendix C).
Since the memory requirements of this approach scale quadratically with the number of tokens, we restrict the maximum document length to $N_{max}=4000$, which is chosen to optimize the utilization of our GPUs. For documents with length $N>N_{max}$, we employ a moving window approach and encode slices of length $N_{max}$ at every stride $S=\frac{N_{max}}{2}$. Relations that span over parts of the document that cannot be captured by any of these windows are discarded and encoded by the mean of the remaining relations. In practice, this affects a fraction of less than $0.2\%$ of the DISRPT instances and thus has a negligible effect on the results.
Note that our approach requires only one forward pass per window encoding multiple relations in parallel and furthermore incorporates the document context into the representation.
We probe both the combined attention scores over all layers as well as the layer-wise representations. [^1] https://github.com/mainlp/discourse_probes We hypothesize that the former will lead to better overall probe accuracy as the understanding of more complex discourse relations typically requires hierarchical processing steps. We include the latter layer-wise analysis to study the structural dynamics of discourse processing in the model.
5 Experimental Setup
Data.
We use the DISRPT benchmark from the DISRPT 2023 shared task Braud et al. (2023), [^2] https://github.com/disrpt/sharedtask2023 a multilingual, multi-domain, and cross-framework (RST, PDTB, SDRT, and DEP) dataset covering $13$ languages from five language families Dryer and Haspelmath (2013), with $224,281$ discourse relation instances from $23$ corpora (details in Appendix B). While all annotations from all frameworks have been represented in a unified format, i.e., a set of discourse unit pairs for which a discourse relation is known to apply Zeldes et al. (2021), the set of discourse relation labels is corpus-specific, preventing multilingual discourse analysis in a directly comparable and generalizable manner. Thus, we map corpus-specific labels in DISRPT to our proposed unified label set.
Models.
We study a wide range of decoder-only LLMs that reflect the diversity of the current state-of-the-art. Generally, we select models with publicly available weights along two dimensions: (1) the size of the model, and (2) the multilinguality of the training corpus. A full list of models including their advertised supported languages and parameter counts is shown in Table 2 in Appendix C.
Figure 2: Mean accuracy over five runs of the probing classifiers trained on the entire DISRPT and full attention representations. The reference system DisCoDisCo achieved a mean accuracy of $47.9\%$ (the red dashed line).
Representing two of the most popular open-weights model families, we include most of the range of Qwen2.5 (Qwen et al., 2025) and Llama3 (Grattafiori et al., 2024) base models. The authors of both models report altered shares of multilingual data in their training corpora, with Qwen2.5 ‘supporting’ $77$% and Llama3 $54$% of the languages included in DISRPT. Covering $62$% and $100$%, we include Mistral-Small-24B (Mistral, 2025) and Emma500 Ji et al. (2024) as further recent multilingual models. BLOOM is another model family that targets multilinguality Scao et al. (2023), covering $54$% of the languages in DISRPT. We include the smaller versions as well as the more recent bloomz-7b1 model trained on additional data. Covering many languages in DISRPT with $85$%, we include the Aya-23 (Aryabumi et al., 2024) and Aya-Expanse (Dang et al., 2024) model families. Finally, we also include Phi-4 Abdin et al. (2024) as a recent monolingual model.
DisCoDisCo.
To better contextualize the performance of our simple probes, we train DisCoDisCo Gessler et al. (2021), the 2021 DISRPT shared task winning system for the discourse relation classification task Zeldes et al. (2021). It was not tested during the 2023 edition, but the reported scores are better than the 2023 winning system on common corpora Braud et al. (2023), justifying its use as a reference system. DisCoDisCo is a Transformer-based model consisting of a sentence-pair classifier. To ensure comparability, we train a single DisCoDisCo model on the entire DISRPT benchmark using xlm-roberta-base Conneau et al. (2020) as a multilingual encoder. We also train dataset-specific models following the setup in Braud et al. (2023). For all scenarios, we report both the dataset-wise accuracy and the mean accuracy.
Figure 3: Mean accuracy over five runs of the Aya-23-35B-probe trained and tested on various partitions of DISRPT.
6 Results & Analysis
6.1 Overall Performance and Comparison
Figure 2 shows the performance of our probes on DISRPT using the unified label set for the $23$ LLMs along with their sizes. [^3] See Table 2 in Appendix C for a detailed overview. Overall, a clear trend emerges wherein larger models generally achieve higher accuracy, but with notable deviations based on their language coverage. Interestingly, even the Llama3-3B probe is able to outperform the fine-tuned DisCoDisCo trained on all languages ($49.1$% vs $47.9$%), while the larger model probes beat it by a margin of up to $10.3$%. Despite their multilingual training, the BLOOM model probes underperform compared to other LLMs and DisCoDisCo. Several factors may be attributed to this including smaller training data, tokenizer effect Toraman et al. (2023), and limited coverage of languages present in DISRPT ($54$%: French, Spanish, Portuguese, English, and Chinese). This aligns with and supplements Dakle et al. (2023)’s findings of the evaluation of BLOOM models on a variety of syntactic and semantic tasks as well as their unsatisfying performance in multilingual settings. In addition, Llama3 exhibits the lowest proportion of DISRPT languages covered, which does not seem to lead to a distinct disadvantage as its probes outperform the similarly sized smaller versions of Qwen2.5 and match its larger counterparts, reporting a higher coverage of ‘supported’ languages.
Examining scaling trends, we observe a log-linear increase in probe performance across both the Llama3 and Qwen2.5 model families. Interestingly, the English-only Phi-4 lags behind the similarly sized Qwen2.5-14B model, suggesting that multilingual training plays a critical role in multilingual discourse analysis. This further reinforces the importance of language coverage in training data, beyond simple model scaling, in achieving robust performance in discourse relation classification. Furthermore, despite being a fine-tuned version of Llama2-7B, Emma500’s multilingual training appears to give it an edge over the more recent and larger Llama3-8B. Similarly, Aya-23-35B’s probe surpasses those of the largest Qwen2.5-70B and Llama3-72B models, despite operating with only half the number of parameters, emphasizing the efficiency of its learned representations. Given this evident edge in the overall probing performance, we focus the rest of our investigation into discourse generalization on Aya-23-35B.
6.2 Language Generalization
We evaluate Aya-23-35B’s performance on the DISRPT test sets across $13$ languages and compare different training conditions: (1) monolingual training (mono-probe), (2) multilingual training with languages from the same language family (multi-lang-probe), and (3) multilingual training using instances from all languages (multi-all-probe). Figure 3 shows the performance of the Aya-23-35B model trained and tested on all and subsets of the data given different training regimes.
multi-all-probe & mono-probe.
The probes trained on the entire DISRPT dataset match or outperform the monolingually trained probes across most test partitions, including the combined test set with all languages. This suggests that multilingual training benefits discourse relation classification, leveraging shared discourse patterns across languages. This is the opposite to what we see in the reference system DisCoDisCo, where the dataset-specific models outperform the model trained on all the data by a margin of $17.8$%. For Farsi and Russian, the best performance is observed in the multi-all-probe, with $+2$% and $+1.2$% over their respective mono-probes. The same generalization is observed for Turkish, Mandarin Chinese, and French, evidenced by an increase in accuracy of $6.9$%, $2$%, and $1.7$% respectively.
(a) languages and language groups
(b) top-level classes
(c) Causal-Argumentative
Figure 4: Layer-wise probe performance by languages and relation classes. Mean accuracy over five runs.
There is one exception to this generalization trend: for the Thai dataset, the mono-probe outperforms multi-all-probe with a margin of $3.2\%$. This is likely due to the fact that the Thai dataset is relatively large and contains only news data and was only annotated with explicit discourse relations Prasertsom et al. (2024), which has been reported to be considerably easier than implicit discourse relations Knaebel (2021); Braud et al. (2023) due to the presence of connectives, which, while not unambiguous, help narrow down the likely senses of relations Webber et al. (2019). The probes for Basque consistently underperform, with the multi-all-probe achieving the same performance as the Basque-only probe. This is likely due to the fact that Basque is considered a language isolate, a language that has no demonstrable genetic relationship with any other language Campbell (2010). Besides, most models do not include it in their list of supported languages (see Table 2 in Appendix). The reference system DisCoDisCo also does not generalize well for Basque: its accuracy drops by $26.4$% compared to the dataset-specific model. Moreover, the mono-probes for Basque and Turkish exhibit lower accuracy, which might mean that models trained on discourse tasks need more language-specific adaptations, particularly for morphologically rich and typologically different languages.
multi-lang-probe.
Training on discourse relation instances from related languages often leads to better generalization than training on a single language. For instance, the multi-lang-probe for the Indo-European languages achieves reasonable performance across the Romance languages in DISRPT: there are improvements over the mono-probe for Portuguese ($+2.5$%), Spanish ($+6.5$%), and Italian ($+3.1$%). For the Germanic and Romance language groups, we observe a similar generalization effect, though to a lesser extent. Notably, the Germanic multi-lang-probe leads to the highest accuracy on the German test set ($+0.8$% over the mono-probe). These are encouraging results as some of these languages do not have a large amount of instances in DISRPT: from the total number of data, Spanish covers $1.7$%, and Italian only $0.7$% (see Table 1 in Appendix B). This suggests that leveraging discourse relation annotations from related language as well as the same underlying framework help with generalization. Adding to that, training on English only, which has the largest number of samples covering more than half of DISRPT, shows surprising generalization effects to other languages, with Portuguese and Spanish achieving an accuracy of $48.2\%$ and $49\%$ respectively, and French reaching $45.3\%$.
6.3 Layer-wise Analysis
The layer-wise probes reveal additional insights into the structural processing of discourse information within the model. Firstly, Figure 4a shows the test set performance by language families. Here, we observe a trend by which the probes in lower to middle layers are the most performant as accuracy improves rapidly in the early layers and stabilizes around layer 1010101015151515. In particular, for Turkish and Thai, the probe accuracy drops by almost $50$% in higher layers. This effect is also observed for the Chinese probe but to a smaller extent. This shows that the discourse representations are best aligned across languages in the middle layers, which could be explained by the findings of Wendler et al. (2024) according to which multilingual computations are carried out in an English-aligned “concept space” in the intermediate layers. Surprisingly, the final layer probes show another increase for Chinese, Turkish, and Thai, which is also reflected in the overall test set accuracy.
For the multi-lang-probes of Romance, Germanic, and Indo-European, the corresponding performance is more constant, closely following the “all” curve (multi-all-probe). Here, the probes of the later layers show a drop of about $5$-$10$% suggesting that their representations’ alignment is more consistent throughout the model. This also indicates that both the higher and lower layers are involved in the discourse processing of LLMs.
Regarding relation classes, Figures 4b shows that structuring and topic-management achieve relatively higher accuracy across layers, peaking in the early-to-mid layers (around $5$-$15$), whereas temporal and causal-argumentative exhibit lower performance. Figure 4c zooms in on the more fine-grained classes: adversative and enablement achieve the highest accuracy overall, with peaks around layers $10$-$20$, while evaluation and explanation remain consistently low. Layer-wise plots for other relation classes are provided in Figure 7 in Appendix D. Overall, different relation classes appear to be encoded at varying layers, with some benefiting from middle-layer representations while others, particularly structure-oriented ones, maintain stable performance across layers. The drop in accuracy in later layers suggests that some discourse information might become less explicitly represented as the model progresses through deeper layers.
6.4 Error Analysis
Figure 5: Confusion Matrix of Labels over multi-all-probe (colors normalized by row).
To better understand the probe performance, we plot a confusion matrix in Figure 5 to study systematic errors. Overall, elaboration is the most frequent label. Here, our probe has a bias of confusing it with labels such as framing and explanation. This suggests that the model struggles to differentiate between content expansion (elaboration) which sometimes overlaps semantically with explanation, and relations that involve additional information but with a primary focus on contextualization (framing). In other words, elaboration is mainly used to give additional information about an entity or a proposition, while framing provides information with a goal to increase a reader’s understanding of an entity or a proposition Mann and Thompson (1988); Carlson and Marcu (2001). This corresponds to the distinction made by Hovy and Maier (1997): elaboration is considered an ideational class that is used to express states of affairs in the world, not including the interlocutors; on the other hand, framing and explanation are considered interpersonal and are expressed by various perlocutionary acts to affect readers’ attitude and beliefs. elaboration is also a highly prevalent relation in general, making it a probable default prediction in ambiguous cases.
In addition, a majority of the comparison instances are predicted as adversative. A qualitative inspection reveals that this is due to some datasets not distinguishing similarities from differences, while the proposed unified label set does. We argue that similarity-based instances highlight commonalities between entities or situations and establish a shared property or behavior, reinforcing coherence by aligning elements. On the other hand, adversative-based cases emphasize differences, opposition, or unexpected alignments, which serve different pragmatic functions. Moreover, similarity-based comparisons often reinforce or extend prior discourse, leading to additive coherence; while adversative-based cases introduce shifts, which require the reader to re-evaluate assumptions and adjust interpretation. By distinguishing these two types, we think models can better capture rhetorical intent and argumentative structure.
7 Conclusion
Our study provides a comprehensive analysis of discourse generalization in LLMs, revealing important patterns in how these models encode discourse structures for cross-lingual transfer. To this end, we first present a unified discourse relation label set, which serves as a foundation for our probing experiments, allowing us to analyze discourse representations beyond individual frameworks. This is also the first work to apply a unified label set across frameworks and languages to multilingual discourse relation classification at scale.
Our probes exhibit generalization, whereby training across languages generally outperforms language–specific probes, which is not the case for the reference system DisCoDisCo. Through multi-faceted analyses, we find that model size alone does not lead to discourse probing success; instead, multilingual training, dataset composition, and language-specific factors play significant roles. While larger models generally perform better, discrepancies such as the under-performing BLOOM and the best-performing Aya-23-35B-probe emphasize the importance of training data quality and architectural optimizations. Surprisingly, our English mono-probe generalizes to Romance languages. Our layer-wise analysis suggests that discourse representations are best aligned across languages in the intermediate layers, with later layers refining these representations for specific relation types. In addition, models struggle with relations requiring implicit reasoning such as explanation.
Overall, our findings highlight the interplay between multilinguality, scaling, and internal representations of LLMs for multilingual discourse processing and provide insights into cross-lingual alignment of discourse relations. Future work can improve cross-lingual alignment by refining representations of challenging relation classes. By advancing our understanding of discourse generalization in LLMs, we contribute towards more interpretable and robust NLP systems capable of nuanced language comprehension.
Limitations
While this work offers a first step towards understanding discourse generalization in LLMs across languages and discourse annotation frameworks, we acknowledge the following limitations related to data, unified representation, and methodological approach.
Firstly, despite being multilingual, the DISRPT benchmark is imbalanced with regard to language coverage. English remains dominant (making up $53.5\%$ of the DISRPT benchmark, as shown in Table 1), which may impact the generalizability of our findings to lower-resource languages. To maintain comparability with prior work, we did not stratify the training data based on language sample sizes. However, we mitigated this by training language-specific probes (mono-probe) and multi-lang-probes based on language families and reporting their respective performance in Figure 3. Similarly, while DISRPT includes multiple domains and genres, some are underrepresented such as conversational data or dialogues. This results in a label imbalance such as topic-adjustment, limiting statistical robustness in such cases. However, this also highlights the need for the creation of more balanced multilingual discourse resources.
Secondly, our approach assumes a unified label set for discourse relations across languages and frameworks, which we present in Section 3. While this enables cross-linguistic and cross-framework discourse analysis, compromises were necessary, which to some extent simplify the complexity and granularity of discourse relations assumed by frameworks such as PDTB and RST. Fully standardizing and harmonizing discourse relations is inherently challenging, and finding a good trade-off between maintaining theoretical assumptions and ensuring practical applicability is crucial yet complex, as evidenced in previous mapping proposals and efforts Benamara and Taboada (2015); Demberg et al. (2019). In addition, segmentation differences exist across frameworks, which can have an impact on the performance of our probes for certain relations such as attribution. This work should thus be interpreted with an awareness of the theoretical and practical difficulties in creating an informed and unified taxonomy suitable for both theoretical studies and computational research. This work should also be viewed as facilitating the development of a unified discourse relation representation for computational discourse modeling such as the effort and initiatives that have been put forward for dependency parsing (Universal Dependencies, de Marneffe et al. 2021), empirical study of anaphora (Universal Anaphora, Poesio et al. 2024), and semantic parsing (Uniform Meaning Representation, Bonn et al. 2024).
Lastly, rather than using highly optimized architectures, we employed relatively simple probing methods, which aligns with our interest in assessing the intrinsic capabilities of LLMs for discourse processing. While achieving state-of-the-art performance was not our primary goal, better performance could likely be achieved by fine-tuning the LLMs. Future work would need to assess the trade-off between generalization and task-specific optimization.
Acknowledgments
We would like to thank the members of the MaiNLP lab for their valuable feedback, especially to Philipp Mondorf, Siyao Peng, and Shijia Zhou. We would also like to express our gratitude to Amir Zeldes for the constructive and valuable feedback. Lastly, we recognize the support for Yang Janet Liu and Barbara Plank through the ERC Consolidator Grant DIALECT 101043235.
Appendix A Unified Label Set: Definitions and Examples
The proposed unified label set adapts the mapping proposal described in Benamara and Taboada (2015) and extends it to be applicable to phenomena frequent in dialogues. In total, there are $17$ labels corresponding to the core discourse relations identified in Bunt and Prasad (2016). Below we provide definitions and list typical framework-specific discourse relations that are mapped to the unified label set.
temporal
is used to map to framework-specific labels that establish a chronological sequence between events or states. RST’s sequence, PDTB’s temporal.asynchronous/synchronous, and SDRT’s temploc and flashback are mapped to this class.
structuring
is mapped to RST-style relations such as list, preparation, disjunction, organization-heading, joint, and textual-organization. SDRT-style relations such as alternation, continuation, and parallel are mapped to this class. PDTB’s expansion.disjunction is mapped to this class.
attribution
informs about the source of information and is considered a discourse relation by RST and SDRT. PDTB’s attribution annotation is considered a separate discourse annotation type and is not included in DISRPT.
comparison
is used to group fine-grained relations that highlight similarities rather than differences between spans or entities such as PDTB’s comparison.similarity and analogy in RST-style corpora.
elaboration
provides additional information about an entity or a proposition. Framework-specific relations that contain elaboration or are prefixed with elab (e.g. elab-process_step, elab-enumember, and Q_Elab) are all mapped to this class. progression used in DEP-style corpora is also mapped to this class.
framing
is used for framework-specific relations that provide a framework for understanding the content of the situation described in the discourse units Benamara and Taboada (2015) such as frame, background, and circumstance
mode
 is used to supply information about how events happens. Commonly mapped fine-grained relations are manner and means in RST-style corpora as well as PDTB’s expansion.manner.
reformulation
corresponds to relations by which one discourse unit re-expresses the meaning of another in a different form and ensure coherence by providing alternative expressions of the same idea. summary and restatement in RST-style corpora and PDTB’s expansion.equivalence are mapped to reformulation.
adversative
highlights incompatibility and covers commonly used discourse relations in all frameworks such as concession and contrast. PDTB’s expansion.exception/Substitution are also mapped to this subclass given their definitions in Webber et al. (2019).
causal
is used to indicate a cause-and-effect relationship. Fine-grained relations that signal that one event, state, or proposition (the cause) leads to or explains another event, state, or proposition (the effect) is mapped to this class. This is one of the most core discourse relation types recognized in all frameworks.
contingency
is used to map condition-based relations such as conditional, unless, unconditional, and contingency.negative-condition.
enablement
is used to connect discourse units where one enables the other. Framework-specific relations such as goal and purpose are mapped to this class.
explanation
is used when the situation described by one argument provides the reason, explanation, or justification for the situation described by the other Webber et al. (2019).
evaluation
is used where one discourse unit provides an assessment, judgment, or commentary on the content of another unit. Framework-dependent relations such as comment and interpretation are mapped to this class.
topic-change
involves a shift or drift in topic that links large textual units Carlson and Marcu (2001). This is used for fine-grained relations that connect multiple, non-contrasting discourse units that are of equal prominence such as joint-other in the eng.rst.gum corpus and PDTB’s expansion.conjunction.
topic-comment
is used for framework-specific relatiosn that involve question-answer pairs or problem-solution pairs, commonly in RST-style and SDRT-style corpora. PDTB’s hypophora is also mapped to this class.
topic-adjustment
is primarily used for cases where a discourse unit modifies, redirects, or adjusts the ongoing topic of discussion such as correction and interrupted, which signal deviations from the expected discourse progression.
Appendix B Data
Table 1 provides an overview of the DISRPT benchmark we use for our experiments. The data produced and used in this paper is in accordance with the original licenses of the underlying resources, as specified in the repository of DISRPT. [^4] https://github.com/disrpt/sharedtask2023
dataset language
language
family
# of relation
instances
% of total
instances
deu.rst.pcc German Indo-European, Germanic 2665 1.19%
eng.dep.covdtb English 4985 2.22%
eng.dep.scidtb 9904 4.42%
eng.pdtb.pdtb 47851 21.34%
eng.pdtb.tedm 529 0.24%
eng.rst.gum 24688 11.01%
eng.rst.rstdt 19778 8.82%
eng.sdrt.stac 12235 5.46%
eus.rst.ert Basque
Language
Isolate
3825 1.71%
fas.rst.prstc Farsi
Indo-European,
Iranian
5191 2.31%
fra.sdrt.annodis French Indo-European, Romance 3338 1.49%
ita.pdtb.luna Italian 1544 0.69%
nld.rst.nldt Dutch
Indo-European,
Germanic
2264 1.01%
por.pdtb.crpc Portuguese Indo-European, Romance 11330 5.05%
por.pdtb.tedm 554 0.25%
por.rst.cstn 4993 2.23%
rus.rst.rrt Russian
Indo-European,
Slavic
34566 15.41%
spa.rst.rststb Spanish Indo-European, Romance 3049 1.36%
spa.rst.sctb 692 0.31%
tha.pdtb.tdtb Thai
Tai-Kadai,
Kam-Tai
10865 4.84%
tur.pdtb.tdb Turkish Altaic, Turkic 3185 1.42%
tur.pdtb.tedm 577 0.26%
zho.dep.scidtb Mandarin Sino-Tibetan, Chinese 1298 0.58%
zho.pdtb.cdtb 5270 2.35%
zho.rst.gcdt 8413 3.75%
zho.rst.sctb 692 0.31%
Table 1: Overview of datasets in DISRPT 2023 for the discourse relation classification task.
Appendix C Models and Probe Training
LLMs.
Table 2 provides an overview of the examined LLMs in this paper, along with their number of parameters, multilingual capabilities, and the proportion of languages in DISRPT covered by the advertised supported languages.
Hyperparameters.
We train our probes using AdamW (Loshchilov and Hutter, 2018) with a batch size of $64$, learning rate of $0.0001$, and weight decay of $0.0001$. To mitigate the class imbalance that is inherent to discourse relation data, we use a weighted cross entropy loss, where each class’ loss is weighted by the inverse square root of the number of samples in the respective class. For the hidden layer, we choose a dimension of $D$=$512$. For the input as well as the hidden layer, we regularize adding a Dropout of $0.2$. Furthermore, we use layer normalization in the hidden layer. For the probes over all model attention scores, depending on the number of samples contained in the train dataset, we train for $60$ epochs and increase that number to ensure at least $10000$ gradient update steps on smaller datasets. For the layer-wise probes, we reduce the number of epochs to $20$ as the probes converge much faster on these smaller representations. Because token-length of encoded sequence quadratically scales the GPU memory required to process the attention matrix, we had to lower the maximum window length for the larger models. Namely, we reduced the maximum window sizes from $N_{max}$=$4000$ to $N_{max}$=$3800$ for Aya-Expanse-32B and to $N_{max}$=$3400$ for Aya-23-35B, Llama3-70B, and Qwen2.5-72B.
Compute.
For each model, we compute one pass over the dataset computing all the attention representations which we cache for the probing experiments. For the smaller models, a cluster equipped with eight Nvidia A100 GPUs was used for around $80$ hours. For the large models of size $70$B and $72$B, we used a cluster equpped with four Nvidia H200 GPUs for about $30$ hours.
Code License.
As specified in the code repository, we release our code under MIT license.
model
name
model
family
# of
params
# languages
supported
deu eng eus fas fra nld por rus spa tur zho tha ita
fraction
supported
Qwen2.5-0.5B Qwen 2.5 0.49B 29 0.769
Qwen2.5-1.5B 1.54B
Qwen2.5-3B 2.77B
Qwen2.5-7B 7.61B
Qwen2.5-14B 14.7B
Qwen2.5-32B 32.5B
Qwen2.5-72B 72.7B
Llama-3.2-1B Llama 3 1.23B 8 0.538
Llama-3.2-3B 3.21B
Llama-3.1-8B 8.03B
Llama-3.1-70B 70.6B
bloom-560m Bloom 0.56B 46
bloom-1b1 1.07B
bloom-3b 3B
bloom-7b1 7.07B
bloomz-7b1 7.07B
aya-expanse-8b Aya 8.03B 23 0.846
aya-expanse-32b 32.3B
aya-23-8B 8.03B
aya-23-35B 35B
phi-4 Phi 14.7B 1 0.077
emma-500-llama2-7b Emma 6.74B 546 1
Mistral-Small-24B-Base-2501 Mistral 23.6B 10 0.615
Table 2: Overview of the included LLMs and their multilingual capabilities for languages in DISRPT.
dataset
Llama-3.1-70B
Llama-3.1-8B
Llama-3.2-1B
Llama-3.2-3B
Mistral-Small-24B-Base-2501
Qwen2.5-0.5B
Qwen2.5-1.5B
Qwen2.5-14B
Qwen2.5-32B
Qwen2.5-3B
Qwen2.5-72B
Qwen2.5-7B
aya-23-35B
aya-23-8B
aya-expanse-32b
aya-expanse-8b
bloom-1b1
bloom-3b
bloom-560m
bloom-7b1
bloomz-7b1
emma-500-llama2-7b
phi-4
DisCoDisCo (all)
DisCoDisCo
deu.rst.pcc 41.1 35.9 31.1 33.5 39.6 24.8 24.2 36.3 37.5 26.6 41.4 30.5 43.3 39.4 38.3 38.8 23.2 27.8 19.4 22.5 23.1 36.2 35.2 25 46.5
eng.dep.covdtb 61.8 56.4 54.7 56.9 57.9 51.1 49.2 60.9 61.6 51.9 62.9 55.6 65.9 58.5 62.5 60.5 49.9 53.8 50.2 53.8 54.7 56.6 58.2 59.9 74.1
eng.dep.scidtb 71.8 63.2 59.5 62.7 64.0 54.4 54.4 66.2 66.8 57.8 67.3 60.5 72.4 63.9 69.5 66.3 54.5 59.1 55.0 58.8 58.9 64.2 63.9 70.7 82.7
eng.pdtb.pdtb 67.4 61.6 56.7 60.6 63.8 51.3 51.7 65.6 67.1 58.5 68.4 61.5 68.8 65.3 67.9 65.3 49.4 56.2 49.4 57.1 56.6 62.0 63.4 59.1 79.9
eng.pdtb.tedm 48.9 53.2 49.0 47.1 52.5 43.8 45.0 52.3 55.7 49.6 53.0 52.6 51.9 58.2 54.6 56.9 40.1 46.4 45.6 47.0 46.4 53.3 52.5 40.7 58.1
eng.rst.gum 54.1 47.4 43.9 45.9 50.5 40.9 40.8 54.4 54.5 45.4 56.3 48.5 55.6 49.7 55.5 52.8 40.9 45.6 40.5 46.5 45.8 49.3 50.2 44.7 64
eng.rst.rstdt 56.0 51.7 51.2 51.7 54.7 48.6 47.6 57.6 57.5 51.1 56.5 53.7 57.6 53.3 55.9 55.2 48.1 51.7 46.7 51.0 50.5 50.1 55.1 49.7 66.1
eng.sdrt.stac 48.0 44.0 41.6 43.9 44.7 40.1 41.7 46.3 46.9 42.7 48.6 42.0 49.1 45.4 47.7 45.2 39.9 41.5 38.7 41.6 41.6 43.1 45.6 45.4 60.4
eus.rst.ert 40.4 29.1 27.4 31.5 35.7 22.9 20.7 36.6 36.8 24.1 38.4 29.6 39.3 30.0 36.5 32.1 27.2 29.4 25.7 30.2 29.4 34.2 34.2 36 62.4
fas.rst.prstc 49.2 43.8 43.1 46.9 47.1 36.7 39.6 49.4 49.0 43.1 51.9 46.8 49.3 48.9 49.5 48.2 37.7 40.2 33.6 40.4 38.6 47.4 46.1 46 52.7
fra.sdrt.annodis 48.4 43.3 41.1 42.5 45.7 41.5 41.9 53.4 52.9 40.9 51.2 47.1 50.7 44.3 47.6 47.6 37.8 42.4 39.3 43.1 42.8 43.2 49.2 26.6 56.2
ita.pdtb.luna 45.7 36.1 32.7 35.4 40.2 18.2 21.1 39.8 41.7 28.7 43.7 34.6 50.9 39.8 44.5 43.9 12.9 23.4 16.6 26.7 24.2 39.9 35.0 47.1 52.1
nld.rst.nldt 40.0 31.9 34.9 33.5 34.5 32.8 33.5 40.7 42.4 33.4 44.9 35.4 41.3 34.9 40.6 35.0 27.4 30.1 25.0 31.3 31.5 34.5 38.0 37.2 56.6
por.pdtb.crpc 65.9 58.8 54.7 57.5 61.5 52.4 50.9 64.9 66.2 54.9 68.0 60.1 67.4 60.7 64.4 63.7 47.0 54.0 50.0 53.6 53.3 63.0 63.8 55.9 75.4
por.pdtb.tedm 57.1 51.6 43.0 47.4 49.3 39.3 41.1 55.2 57.0 45.1 58.7 49.6 61.4 53.2 57.4 55.4 39.7 43.9 39.8 46.7 45.7 50.4 54.1 49.2 66.2
por.rst.cstn 63.9 57.9 55.6 57.8 59.7 55.6 52.9 62.4 61.9 54.8 62.5 57.2 65.3 58.9 62.2 59.9 49.0 55.2 49.1 56.5 54.8 57.2 59.0 53.7 68.8
rus.rst.rrt 56.0 52.4 50.4 51.3 53.9 47.6 47.1 57.2 57.9 50.6 59.2 53.7 58.6 55.7 58.5 55.3 41.8 45.1 42.9 45.2 42.3 54.6 54.5 50.6 64.7
spa.rst.rststb 48.4 40.0 39.7 39.5 46.9 35.2 34.9 44.6 47.7 39.2 47.6 43.7 52.9 47.5 50.8 48.4 31.7 38.0 31.4 40.4 38.7 46.4 40.9 42 61
spa.rst.sctb 62.9 66.9 63.5 63.5 65.0 60.4 56.2 60.9 61.9 62.0 68.3 63.1 69.7 64.2 68.9 66.3 51.2 55.7 50.6 55.5 56.0 64.3 59.6 46.5 66
tha.pdtb.tdtb 85.2 76.3 65.0 72.1 77.6 55.2 56.7 81.5 82.1 69.3 83.4 72.7 83.4 75.7 81.9 76.4 27.3 37.1 25.1 39.0 34.4 77.6 78.4 90.6 84.3
tur.pdtb.tdb 52.8 47.7 42.8 44.5 48.3 39.2 38.0 54.2 56.5 43.6 56.5 48.5 57.9 51.7 56.3 49.4 33.2 34.8 30.5 36.1 32.6 50.8 50.2 40.3 68.7
tur.pdtb.tedm 48.2 43.2 39.2 43.0 48.1 32.9 32.2 49.6 50.2 35.9 52.6 43.4 49.7 44.2 50.2 46.3 17.9 24.2 20.5 23.8 23.7 48.1 42.7 27.5 53.3
zho.dep.scidtb 60.7 48.3 36.3 51.6 48.4 32.8 30.8 49.8 53.9 39.4 52.8 42.0 61.6 48.0 56.8 53.8 33.9 37.4 32.9 43.3 42.0 46.7 47.3 53.5 72.1
zho.pdtb.cdtb 75.9 71.2 70.6 72.6 75.4 62.7 65.4 76.3 77.1 66.2 78.0 71.8 76.3 74.5 78.4 74.3 57.2 68.7 55.8 66.3 66.8 65.6 73.7 68.3 93.8
zho.rst.gcdt 52.4 46.1 41.6 43.1 48.5 35.5 35.6 54.9 53.4 44.1 53.9 47.9 56.4 50.0 54.9 51.5 36.0 41.6 37.5 43.8 43.0 51.9 49.5 45.5 64.3
zho.rst.sctb 57.5 45.4 33.5 40.8 44.8 35.2 38.0 44.8 49.2 35.2 54.0 41.9 56.1 38.9 51.9 50.1 41.5 35.8 38.9 39.1 36.9 42.5 40.6 34 56.6
average 56.1 50.1 46.3 49.1 52.2 42.0 42.0 54.5 55.6 45.9 56.9 49.8 58.2 52.1 56.3 53.8 38.3 43.0 38.1 43.8 42.9 51.3 51.6 47.9 65.7
Table 3: Dataset accuracy scores by LLM probe averaged over five runs. Last two columns refer to the DisCoDisCo reference system trained on all languages (all) and trained a multiple models with different encoders per dataset.
Figure 6: Sum of the absolute dot product between the weight matrices of the probe by layer.
Appendix D Additional Results and Analysis
Table 3 provides dataset accuracy scores for all examined LLM probes as well as the reference system DisCoDisCo’s performance.
Figure 6 shows the layer-wise sums of the dot-products of the weight matrices. The magnitude of these scores can be interpreted as a feature importance and confirm that earlier layers play a crucial role in predictions, although the higher layers also show higher scores from layers $30$ to $38$, indicating a progressive refinement of discourse representations. This is in accordance with previous literature, which has shown that the lower layers of LLMs are mostly involved with encoding, while the upper layers are involved in generating and updating the next token prediction Ju et al. (2024).
(a) thematic
(b) topic-management
Figure 7: Layer-wise probe performance by relation classes. Mean accuracy over five runs.
Appendix E Use of AI Assistants
The implementation of this work has been written with the support of code completions of an AI coding assistant, namely GitHub Copilot. Completions were mostly single lines up to a few lines of code and were always checked carefully to ensure their functionality and safety. Furthermore, we did our best to avoid accepting code completions that would be incompatible with the license of our code or could be regarded as plagiarism. We also include this statement in the README.md of the codebase.

### v2

@@
 Lastly, we discuss challenges and biases in LLM discourse representations, providing insights into the limitations and potential avenues for improving discourse modeling in multilingual and generalization settings.
+The implementations of our experiments are available on GitHub.
+[^1] https://github.com/mainlp/discourse_probes
 2 Related Work
@@
 (2023).
-A similar effort to unify discourse relations has been made by the 2021 DISRPT Shared Task Zeldes et al.
-(2021) which introduced the first iteration of the discourse relation classification task in a unified format.
+In particular, while Bunt and Prasad (2016) identified a set of core discourse relations, it did not cover DEP and was limited to English and French only.
+The examined corpora in their work also did not cover discourse phenomena concerning pragmatics or textual organization, both of which are indispensable aspects in discourse analysis.
+For instance, background and motivation are two RST-style relations that are not present in the examined RST Discourse Treebank (RST-DT, Carlson et al.
+2003).
+Both relations are expressed by various perlocutionary acts to affect readers’ or speakers’ attitude and beliefs.
+To address these limitations, we first conduct an extensive review of previous work on discourse relation mapping proposals and present a unified label set (§3) to enable empirical studies in a multilingual setting, which has not been systematically explored before.
+The 2021 DISRPT Shared Task Zeldes et al.
+(2021) introduced the first iteration of the discourse relation classification task in a unified format.
 It leveraged shared foundational assumptions across frameworks.
 However, no unified discourse relation labels were proposed, meaning that each dataset has its own label set, even for the ones that come from the same framework.
-To address this, we first conduct an extensive review of previous work on discourse relation mapping proposals and build upon existing work to present a unified label set, which we describe in Section 3.
+We thus leverage this resource and propose a unified label set that is the first to be empirically tested in discourse relation classification across $13$ languages, four frameworks, and $26$ datasets, which cover various, modern genres, domains, and modalities.
 Probing for Linguistic Representation and Generalization.
@@
 (2025).
-Previous work has examined the ability of PLMs/LLMs to understand discourse such as coreference resolution Gan et al.
+Previous work has examined the ability of PLMs/LLMs to understand discourse Gan et al.
 (2024); Saputa et al.
+(2024); Miao et al.
 (2024), but their investigations are either limited to framework-dependent representations, monolingual datasets, or focus on single domains.
@@
 3 A Unified Label Set
-Building up on previous work and effort on mapping discourse relations across corpora and frameworks Benamara and Taboada (2015); Bunt and Prasad (2016); Liu and Zeldes (2023), we present a unified set of $17$ discourse relation classes that are identified as core discourse relations by Bunt and Prasad (2016) and are attested in various domains and text types.
+Building up on previous effort on mapping discourse relations across corpora and frameworks Benamara and Taboada (2015); Bunt and Prasad (2016); Liu and Zeldes (2023), we present a unified set of $17$ discourse relation classes to facilitate empirical investigation that is not constrained by framework-dependent discourse representations.
 This unified label set is motivated by both theoretical groundings and empirical studies, and takes annotation guidelines into considerations.
@@
 (2016).
-We describe the top-level classes below and include definitions and examples for all $17$ relation labels in Appendix A.
+Through a series of experiments, we demonstrate how our unified label set generalizes across languages and frameworks, providing a foundation for future empirical studies.
+Below we describe the top-level classes and include definitions and examples for all $17$ relation labels in Appendix A.
 temporal is mapped to framework-specific labels that establish a chronological sequence between events or states.
@@
 RST’s sequence, PDTB’s Temporal.Asynchronous/Synchronous, and SDRT’s temploc and flashback (following Muller et al.
-(2012)) all fall under this class.
+2012) all fall under this class.
 structuring corresponds to fine-grained discourse relations that organize the structure of a text or conversation without necessarily conveying content-based meaning, connect discourse units of distinct context and equal prominence, and help guide the reader or listener through the discourse.
@@
 RST’s summary and restatement and PDTB’s expansion.equivalence are mapped to reformulation.
+Figure 2: Schematic visualization of how the attention representations are obtained.
+We probe the concatenation of pooled representations $(\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C})$.
+Scores mapping from or to context-tokens are not considered.
+Note that we study decoder-only models where attention maps only to previous tokens.
 causal-argumentative contains subclasses that indicate a causal relation or involve rhetorical reasoning that shapes the coherence and persuasiveness of an argument: causal, adversative, explanation, evaluation, contingency, and enablement.
-adversative is defined as connecting discourse units for which some incompatibility is being highlighted and covers commonly used discourse relations in virtually all frameworks such as concession and contrast.
+adversative is defined as connecting discourse units for which some incompatibility is being highlighted and covers commonly used discourse relations such as concession and contrast.
 topic-management contains three subclasses that cover discourse phenomena that involve interaction, topic shifts, and question-answer pairs: topic-adjustment, topic-change, and topic-comment.
@@
 In the computational graph of the Transformer, attention layers provide the only connections between the next token prediction and previous token positions of the input sequence.
-We therefore hypothesize that generating predictions relying on discourse-level information necessarily has to involve the connections provided by the self-attention layers.
+We therefore argue that generating predictions relying on discourse-level information necessarily has to involve the connections provided by the self-attention layers.
 To uncover the extent to which discourse information is represented, we thus propose to probe the attention scores between the tokens contained in the two input sequences for discourse relation classification.
+Note that we do not assume that the processing of discourse-level phenomena is exclusively located in the attention layers.
+Rather, observing attention scores provides a lightweight and scalable approach to our research question.
 Probing requires a fixed length representation to enable training a classifier.
@@
 (2021), we use maximum pooling to turn the attention scores of variable token-length sentences into a fixed-length relational representation.
-To be more precise, for a document $d$ of length $N$ tokens, we propose to compute the full attention matrix $\mathbf{X}\in\mathbb{R}^{L\times H\times N\times N}$ where $L$ is the number of layers and $H$ is the number of attention heads in each layer.
+Our approach is illustrated in Figure 2.
+To be more precise, for a document $d$ of length $N$ tokens, we propose to compute the full attention matrix $\mathbf{X}\in\mathbb{R}^{L\times H\times N\times N}$ of attention scores where $L$ is the number of layers and $H$ is the number of attention heads in each layer.
+We do this by inputting the whole document $d$ into the model at once computing a single forward pass.
+Since we are only interested in the model internals, we ignore the outputs and thus do not impose any specific decoding strategy.
 Let $I_{1}=(i_{1},i_{2},...,i_{N_{1}})$, $I_{2}=(j_{1},j_{2},...,j_{N_{2}})$ be the token indices of two sequences $s_{1},s_{2}\subset d$ where $I_{1}<I_{2}$.
@@
 $\begin{split}\mathbf{C}&=\left(\max(\{\mathbf{X}_{i,j,k,l}|k\in I_{1},l\in I_{2}\})\right)_{i,j}\\ \mathbf{D}_{m}&=\left(\max(\{\mathbf{X}_{i,j,k,l}|k\in I_{m},l\in I_{m}\})\right)_{i,j}\\ \end{split}$ (1)
-In other words, $\mathbf{C},\mathbf{D}_{1},\mathbf{D}_{2}$ are the matrices of pooled attention scores between and within the two spans respectively (we only consider the lower half of the attention matrix, as for decoder models the upper half is masked).
-We probe on the flattened concatenation $\mathbf{a}\in\mathbb{R}^{3LH}$ of these representations.
-Here, we follow Tenney et al.
-(2019) and use a two-layer MLP with $\mathrm{tanh}$ and Sigmoid activations $\sigma$.
-Our probe is thus a classifier of the form
+In other words, $\mathbf{C},\mathbf{D}_{1},\mathbf{D}_{2}$ are the matrices of maximum-pooled attention scores between and within the two spans respectively (we only consider the lower half of the attention matrix, as for decoder models the upper half is masked).
+We also ablate other strategies such as mean pooling and using subsets of the attention scores described (see Table 4 and Table 5 in Appendix D), and find that the proposed setup strikes the best balance between performance and size of the representations.
+We probe the flattened concatenation
+$\mathbf{a}=flatten((\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}))\in\mathbb{R}^{3LH}$ (2)
+Here, we follow Tenney et al. (2019) and use a two-layer MLP with $\mathrm{tanh}$ and Sigmoid activations $\sigma$. Our probe is thus a classifier of the form
 $\mathbf{y}=\sigma(\mathbf{W}_{2}\mathrm{tanh}(\mathbf{W}_{1}\mathbf{a}+\mathbf{b}_{1})+\mathbf{b}_{2})$ (3)
@@
 We probe both the combined attention scores over all layers as well as the layer-wise representations.
-[^1] https://github.com/mainlp/discourse_probes We hypothesize that the former will lead to better overall probe accuracy as the understanding of more complex discourse relations typically requires hierarchical processing steps.
+We hypothesize that the former will lead to better overall probe accuracy as the understanding of more complex discourse relations typically requires hierarchical processing steps.
 We include the latter layer-wise analysis to study the structural dynamics of discourse processing in the model.
@@
 For all scenarios, we report both the dataset-wise accuracy and the mean accuracy.
-Figure 3: Mean accuracy over five runs of the Aya-23-35B-probe trained and tested on various partitions of DISRPT.
 6 Results & Analysis
 6.1 Overall Performance and Comparison
-Figure 2 shows the performance of our probes on DISRPT using the unified label set for the $23$ LLMs along with their sizes.
-[^3] See Table 2 in Appendix C for a detailed overview.
+Figure 3 shows the performance of our probes on DISRPT using the unified label set for the $23$ LLMs along with their sizes (see Table 2 in Appendix C for a detailed overview).
 Overall, a clear trend emerges wherein larger models generally achieve higher accuracy, but with notable deviations based on their language coverage.
@@
 (2023), and limited coverage of languages present in DISRPT ($54$%: French, Spanish, Portuguese, English, and Chinese).
-This aligns with and supplements Dakle et al.
-(2023)’s findings of the evaluation of BLOOM models on a variety of syntactic and semantic tasks as well as their unsatisfying performance in multilingual settings.
-In addition, Llama3 exhibits the lowest proportion of DISRPT languages covered, which does not seem to lead to a distinct disadvantage as its probes outperform the similarly sized smaller versions of Qwen2.5 and match its larger counterparts, reporting a higher coverage of ‘supported’ languages.
+This is consistent and further supplements Dakle et al.
+(2023)’s findings on the evaluation of BLOOM models across a variety of syntactic and semantic tasks as well as their unsatisfying performance in multilingual settings.
+In addition, Llama3 exhibits the lowest proportion of DISRPT languages covered, which does not seem to lead to a disadvantage as its probes outperform the similarly sized smaller versions of Qwen2.5 and match its larger counterparts, reporting a higher coverage of ‘supported’ languages.
 Examining scaling trends, we observe a log-linear increase in probe performance across both the Llama3 and Qwen2.5 model families.
@@
 Given this evident edge in the overall probing performance, we focus the rest of our investigation into discourse generalization on Aya-23-35B.
+Figure 4: Mean accuracy over five runs of the Aya-23-35B-probe trained and tested on various partitions of DISRPT.
 6.2 Language Generalization
@@
 This is likely due to the fact that Basque is considered a language isolate, a language that has no demonstrable genetic relationship with any other language Campbell (2010).
-Besides, most models do not include it in their list of supported languages (see Table 2 in Appendix).
+Besides, most models do not include it in their list of supported languages (see Table 2 in Appendix C).
 The reference system DisCoDisCo also does not generalize well for Basque: its accuracy drops by $26.4$% compared to the dataset-specific model.
@@
 Notably, the Germanic multi-lang-probe leads to the highest accuracy on the German test set ($+0.8$% over the mono-probe).
-These are encouraging results as some of these languages do not have a large amount of instances in DISRPT: from the total number of data, Spanish covers $1.7$%, and Italian only $0.7$% (see Table 1 in Appendix B).
+These are encouraging results as some of these languages do not have a large amount of instances in DISRPT: from the total number of data, Spanish covers $1.7$%, German $1.19$%, and Italian only $0.7$% (see Table 1 in Appendix B).
 This suggests that leveraging discourse relation annotations from related language as well as the same underlying framework help with generalization.
-Adding to that, training on English only, which has the largest number of samples covering more than half of DISRPT, shows surprising generalization effects to other languages, with Portuguese and Spanish achieving an accuracy of $48.2\%$ and $49\%$ respectively, and French reaching $45.3\%$.
+Adding to that, training on English only, which has the largest number of samples covering more than half of DISRPT, shows surprising generalization to other languages, with Portuguese and Spanish achieving an accuracy of $48.2\%$ and $49\%$ respectively, and French reaching $45.3\%$.
 6.3 Layer-wise Analysis
@@
 (2024) according to which multilingual computations are carried out in an English-aligned “concept space” in the intermediate layers.
+This result is also aligned with concurrent work by Skean et al.
+(2025) which finds that final-layer representations are consistently outperformed by intermediate representations across a range of tasks and architectures.
+Our work supplements this by extending their finding to linguistic tasks such as discourse relation classification.
 Surprisingly, the final layer probes show another increase for Chinese, Turkish, and Thai, which is also reflected in the overall test set accuracy.
+This could indicate that their probes benefit from language-specific features, which might be more prevalent in layers close to the outputs if we assume a shared, compressed representation space in intermediate layers as proposed by Wendler et al.
+(2024) and Deletang et al.
+(2024).
 For the multi-lang-probes of Romance, Germanic, and Indo-European, the corresponding performance is more constant, closely following the “all” curve (multi-all-probe).
@@
 6.4 Error Analysis
-Figure 5: Confusion Matrix of Labels over multi-all-probe (colors normalized by row).
 To better understand the probe performance, we plot a confusion matrix in Figure 6 to study systematic errors.
@@
 This suggests that the model struggles to differentiate between content expansion (elaboration) which sometimes overlaps semantically with explanation, and relations that involve additional information but with a primary focus on contextualization (framing).
+Figure 6: Confusion matrix of labels over multi-all-probe (colors normalized by row).
 In other words, elaboration is mainly used to give additional information about an entity or a proposition, while framing provides information with a goal to increase a reader’s understanding of an entity or a proposition Mann and Thompson (1988); Carlson and Marcu (2001).
@@
 While larger models generally perform better, discrepancies such as the under-performing BLOOM and the best-performing Aya-23-35B-probe emphasize the importance of training data quality and architectural optimizations.
-Surprisingly, our English mono-probe generalizes to Romance languages.
-Our layer-wise analysis suggests that discourse representations are best aligned across languages in the intermediate layers, with later layers refining these representations for specific relation types.
+Surprisingly, we find that some of our probes generalize to languages unseen during probe training.
+For example, the English mono-probe generalizes to Romance languages.
+Furthermore, our layer-wise analysis suggests that discourse representations are best aligned across languages in the intermediate layers, with later layers refining these representations for specific relation types.
 In addition, models struggle with relations requiring implicit reasoning such as explanation.
 Overall, our findings highlight the interplay between multilinguality, scaling, and internal representations of LLMs for multilingual discourse processing and provide insights into cross-lingual alignment of discourse relations.
+Understanding which discourse relations are well-captured by LLMs and which are not could help improve discourse parsing models by highlighting gaps in current representation.
+In addition, our findings may be useful for designing better discourse-aware pre-training or fine-tuning strategies.
 Future work can improve cross-lingual alignment by refining representations of challenging relation classes.
+More generally, our method provides a systematic way to investigate discourse representation encoded in LLMs, making it a useful tool for answering linguistic questions that can be formulated into a consecutive span representation.
 By advancing our understanding of discourse generalization in LLMs, we contribute towards more interpretable and robust NLP systems capable of nuanced language comprehension.
@@
 However, this also highlights the need for the creation of more balanced multilingual discourse resources.
-Secondly, our approach assumes a unified label set for discourse relations across languages and frameworks, which we present in Section 3.
+Secondly, our approach assumes a unified label set for discourse relations across languages and frameworks, which we present in §3.
 While this enables cross-linguistic and cross-framework discourse analysis, compromises were necessary, which to some extent simplify the complexity and granularity of discourse relations assumed by frameworks such as PDTB and RST.
@@
 We would also like to express our gratitude to Amir Zeldes for the constructive and valuable feedback.
+We thank the reviewers and the meta-reviewer for their feedback and suggestions.
 Lastly, we recognize the support for Yang Janet Liu and Barbara Plank through the ERC Consolidator Grant DIALECT 101043235.
@@
 attribution
-informs about the source of information and is considered a discourse relation by RST and SDRT.
+informs about the source of information, which is useful and crucial for many real-world applications such as misinformation detection and fact-checking.
+It is considered a discourse relation by RST, SDRT, and DEP.
 PDTB’s attribution annotation is considered a separate discourse annotation type and is not included in DISRPT.
+It is worth pointing out that attribution is not considered a core discourse relation by Bunt and Prasad (2016), but it is one of the most frequent discourse relations in most frameworks.
 comparison
@@
 For the smaller models, a cluster equipped with eight Nvidia A100 GPUs was used for around $80$ hours.
-For the large models of size $70$B and $72$B, we used a cluster equpped with four Nvidia H200 GPUs for about $30$ hours.
+For the large models of size $70$B and $72$B, we used a cluster equipped with four Nvidia H200 GPUs for about $30$ hours.
 Code License.
-As specified in the code repository, we release our code under MIT license.
+As specified in the code repository, [^4] https://github.com/mainlp/discourse_probes we release our code under MIT license.
 model
@@
 Mistral-Small-24B-Base-2501 Mistral 23.6B 10 0.615
-Table 2: Overview of the included LLMs and their multilingual capabilities for languages in DISRPT.
+Table 2: Overview of the included LLMs and their multilingual capabilities for languages in DISRPT 2023.
+Figure 7: Sum of the absolute dot product between the weight matrices of the probe by layer.
+Appendix D Additional Results and Analysis
+Table 3 provides dataset accuracy scores for all examined LLM probes as well as the reference system DisCoDisCo’s performance.
+Figure 7 shows the layer-wise sums of the dot-products of the weight matrices.
+The magnitude of these scores can be interpreted as a feature importance and confirm that earlier layers play a crucial role in predictions, although the higher layers also show higher scores from layers $30$ to $38$, indicating a progressive refinement of discourse representations.
 dataset
@@
 Table 3: Dataset accuracy scores by LLM probe averaged over five runs.
-Last two columns refer to the DisCoDisCo reference system trained on all languages (all) and trained a multiple models with different encoders per dataset.
-Figure 6: Sum of the absolute dot product between the weight matrices of the probe by layer.
-Appendix D Additional Results and Analysis
-Table 3 provides dataset accuracy scores for all examined LLM probes as well as the reference system DisCoDisCo’s performance.
-Figure 6 shows the layer-wise sums of the dot-products of the weight matrices.
-The magnitude of these scores can be interpreted as a feature importance and confirm that earlier layers play a crucial role in predictions, although the higher layers also show higher scores from layers $30$ to $38$, indicating a progressive refinement of discourse representations.
-This is in accordance with previous literature, which has shown that the lower layers of LLMs are mostly involved with encoding, while the upper layers are involved in generating and updating the next token prediction Ju et al.
-(2024).
+The last two columns refer to the DisCoDisCo reference system trained on all languages (all) and trained a multiple models with different encoders per dataset.
+Llama-3.1-70B
+Llama-3.1-8B
+Llama-3.2-1B
+Llama-3.2-3B
+Mistral-Small-24B-Base-2501
+Qwen2.5-0.5B
+Qwen2.5-1.5B
+Qwen2.5-14B
+Qwen2.5-32B
+Qwen2.5-3B
+Qwen2.5-72B
+Qwen2.5-7B
+aya-23-35B
+aya-23-8B
+aya-expanse-32b
+aya-expanse-8b
+bloom-1b1
+bloom-3b
+bloom-560m
+bloom-7b1
+bloomz-7b1
+emma-500-llama2-7b
+phi-4
+$\mathbf{C}$ (inter) 53.8 45.5 41.5 43.4 45.5 35.1 35.6 49.1 52.9 40.8 55.8 42.4 53.7 47.6 54.3 47.0 32.4 39.9 32.3 40.3 40.1 44.4 46.9
+$\mathbf{D}_{1},\mathbf{D}_{2}$ (intra) 46.0 38.2 35.2 37.3 40.1 33.3 32.9 43.5 44.5 33.4 47.4 38.3 45.5 36.0 39.9 40.3 28.5 30.8 29.0 31.9 31.3 40.4 41.3
+$\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}$ (all) 55.9 50.8 46.2 48.6 52.5 41.6 42.2 54.2 55.8 46.0 56.8 50.3 58.2 52.6 56.0 54.0 37.8 42.6 37.0 43.5 43.9 51.4 51.9
+Table 4: Overall accuracy scores by LLM probe averaged over five runs training on the entire DISRPT (probe representation ablation).
+We ablate different types of attention representations used for probing and find that using all described matrices $\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}$ yields the best results.
+Llama-3.1-70B
+Llama-3.1-8B
+Llama-3.2-1B
+Llama-3.2-3B
+Mistral-Small-24B-Base-2501
+Qwen2.5-0.5B
+Qwen2.5-1.5B
+Qwen2.5-14B
+Qwen2.5-32B
+Qwen2.5-3B
+Qwen2.5-72B
+Qwen2.5-7B
+aya-23-35B
+aya-23-8B
+aya-expanse-32b
+aya-expanse-8b
+bloom-1b1
+bloom-3b
+bloom-560m
+bloom-7b1
+bloomz-7b1
+emma-500-llama2-7b
+phi-4
+mean 55.6 48.7 45.2 47.8 48.6 39.4 40.5 50.5 51.7 42.9 54.4 46.4 54.2 48.6 53.1 50.1 35.1 40.0 34.6 40.1 39.9 49.7 48.3
+max 55.9 50.8 46.2 48.6 52.5 41.6 42.2 54.2 55.8 46.0 56.8 50.3 58.2 52.6 56.0 54.0 37.8 42.6 37.0 43.5 43.9 51.4 51.9
+mean,max 56.1 51.0 46.5 48.6 52.8 42.6 42.8 54.0 56.0 47.1 56.1 50.8 58.4 52.8 56.9 54.2 38.5 43.7 38.4 43.2 44.0 51.5 52.0
+Table 5: Overall accuracy scores by LLM probe averaged over five runs training on the entire DISRPT (pooling strategy ablation).
+We ablate different attention-head-wise pooling strategies, namely mean pooling, maximum pooling, and the concatenation of both on the attention score matrices $\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{C}$.
+We find that concatenating both generally yields the best results, though usually only by a margin of less than $1\%$.
+To keep the size of the representations shorter, we thus opt to only keep the maximum pooling.
 (a) thematic